On January 1, X9, Gerald received his 50 percent profits and capital interest in High Air, LLC, in exchange for $2,700 in cash and real property with a $3,700 tax basis secured by a $2,700 nonrecourse mortgage. High Air reported a $15,700 loss for its X9 calendar year. How much loss can Gerald deduct, and how much loss must he suspend if he only applies the tax basis loss limitation?
Answer #1

Gerald deduct can deduct $6050 and he must suspend $1800 if he only applies the tax basis loss limitation.

Gerald's initial tax basis =2700+4700-2700+(50%*2700)) = $6050

Gerald would be allocated 50% = 50%*15700 = $7850

Gerald could deduct = $6050

Remaining loss would be suspended and carried forward indefinitely = 7850-6050 = $1800